OpenStudy (anonymous):

Find two consecutive even integers such that the sum of the second and fourth is 76.

OpenStudy (anonymous):

Our even numbers are\[\{a,a+2,a+4,a+6,...,a+2*n\}\]We have to take the second and fourth\[second:~a+2\\fourth:~a+6\]The sum of them will give us the value of "a"\[(a+2)+(a+6)=76\]\[2a+8=76\]\[\boxed{a=34}\]Replacing this new value\[second:~34+2=36\\fourth:~34+6=40\]